### v1

# On edge-colouring-games by Erdős, and Bensmail and Mc Inerney

Abstract
We consider two games proposed by Erdős, and one game by Bensmail and Mc Inerney, all with the same setup of two players alternately colouring one edge of a clique. We give observations and particular behaviour for each of these problems, and prove a first reduction towards confirming the conjecture by Bensmail and Mc Inerney. We state a conjecture for Erdős’ game on the largest induced maximum degree, and extensions to edge-transitive and, respectively, regular graphs.
1 Introduction
We consider three games: two proposed by Erdős [1] (popularised last year at BCC by Cameron [4] and as problem in the database of Erdős’ problems, https://www.erdosproblems.com/778), and one by Bensmail and Mc Inerney [2], all with a similar setup.
Consider a graph $G$ of order $n$ and size $m.$ Alice and Bob play a game on the edges of $G$, taking turns to colour them red (Alice) and blue (Bob). Let $A$ be the set of red edges and $B$ the set of blue edges at the end. When all edges are coloured, a score is assigned. Here Alice will have $a$ points and Bob have $b$ points, where $a$ and $b$ depend on the final colouring of $G$, i.e., on $G[A]$ and $G[B]$ (the graphs induced by the edges in their colour respectively). Here, $a$ and $b$ will be bounded by the order of the graph, $n$. Both players play optimally, ensuring that their score is as high as possible relative to the other’s score. The score of the game, $s(G)$, equals the difference $a-b$. Alice tries to maximise $s(G)$ and Bob tries to minimise $s(G)$. To ensure a unique outcome $(a,b)$, we assume Alice tries to maximise $a$ for the optimal value of $s=a-b$, and Bob tries to minimise it for the optimal value of $s=a-b$. Equivalently, considering the outcome-function $f(G)=(C+1)a-Cb=C(a-b)+a$ for some $C\geq n$, Alice tries to make $f(G)$ as large as possible and Bob wants $f(G)$ to be as small as possible.
By Zermelo’s theorem [8], we know $f(G)$ and $s(G)$ are uniquely determined when both players play optimally. So, there is a unique optimal outcome $(a,b).$
The three games are called the Clique game, the Star game and the vertex-capturing game.
  • In the Clique game, $\operatorname{Clique}$, $a=\omega(G[A])$ and $b=\omega(G[B])$.
  • In the Star game, $\operatorname{Star}$, $a=\Delta(G[A])$ and $b=\Delta(G[B])$.
  • In the vertex-capturing game, $\operatorname{VC}$, at the end of the game; $a$ is the number of vertices for which most incident edges are red, and $b$ is the number of vertices for which most incident edges are blue ($a+b<n$ is possible if there are vertices with half of their incident edges in either colour).
Each game played on a graph $G$ is denoted by $\operatorname{Clique}(G),\operatorname{Star}(G)$ or $\operatorname{VC}(G)$ respectively, when $G=K_{n}$, we write $\operatorname{Clique}(n),\operatorname{Star}(n)$ or $\operatorname{VC}(n).$
The outcome $(a,b)$ and score $a-b$ of the games will be denoted by respectively $\operatorname{out}_{\omega},\operatorname{out}_{\Delta},\operatorname{out}_{\operatorname{VC}}$ and $s_{\omega},s_{\Delta},s_{\operatorname{VC}}$.
By a strategy-stealing argument, one can prove that the first player has an advantage (or at least will never perform worse than the second player). For details, one can read the proof of [2, Thm. 2.1 & 2.2], which applies to all three games.
Proposition 1.
For every graph $G$, $s_{\omega}(G),s_{\Delta}(G)\in\{0,1\}$ and $s_{\operatorname{VC}}(G)\in\{0,1,2\}$.
For the Clique- and Star-game, we say a graph $G$ is a player 1 win, if $s_{\omega}(G),s_{\Delta}(G)=1$ (the maximum score possible for player $1$), and otherwise it is a player $2$ win.
Erdős conjectured that for $\operatorname{Clique}(n)$, the second player can always force a win as long as $n\not=2,$ and asked the same question for the Star game.
Conjecture 2 ([1]).
$\operatorname{Clique}(n)$ is a player $2$ win, $s_{\omega}(n)=0$, for all $n\in\mathbb{N}\setminus\{2\}.$
Question 3 ([1]).
For every $n\in\mathbb{N},$ which player wins the game $\operatorname{Star}(n)$ (when is $s_{\Delta}(n)=1$)?
Bensmail and Mc Inerney [2, Conj. 7.2] conjectured the exact outcome for the vertex-capturing game (on cliques), and confirmed it for $n\leq 6$, in [2, Prop. 7.6].
Conjecture 4 ([2]).
$s_{\operatorname{VC}}(n)=\begin{cases}2\mbox{ if }n=2\\ 1\mbox{ if }n\equiv 3\pmod{4}\\ 0\mbox{ otherwise. }\\ \end{cases}$
Despite the interest of many mathematicians, little progress has been made. In our work, we highlight the underlying difficulties that may explain why it remains unresolved.
In Section 2, we consider the three games for Colex graphs. A Colex graph $\mathcal{C}(m)$, where $m=\binom{m_{1}}{2}+m_{2}$ with $0\leq m_{2}<m_{1}$, is the graph formed by adding $m_{2}$ edges between $K_{m_{1}}$ and an additional vertex (if $m_{2}>0$). See Fig. 1 for examples where $(m_{1},m_{2})\in\{(4,3),(5,2)\}.$ As a class of graphs that smoothly extend cliques, this serves as a showcase to challenge certain intuitive expectations by exhibiting particular behaviours. The results are determined by computer code (which implements the idea of Zermelo’s theorem [8]), explained in Section 4.
Having solved some small cases, in Section 3, we extend an observation by [6] using strategy stealing, confirming that $\operatorname{Clique}(n)$ being a win for the first player implies that $\operatorname{Clique}(n+3)$ is a win for the second player. We also prove one implication between the validity of 4 for two residue classes modulo $4$. [^1] This proof by strategy stealing is similar to reductions in [6] and was obtained before their publication, as communicated with them.
Proposition 5.
If 4 is true for every $n\equiv 3\pmod{4},$ then it is true for every $n\equiv 1\pmod{4}$ as well.
Finally, we give our intuition, formulated in conjectures for the Clique and Star game (stated in the framework of edge-transitive and regular graphs). In particular, we are the first who dare to conjecture the answer for 3.
Conjecture 6.
With the exception of $K_{2}$, for every edge-transitive graph $G$, the Clique game is winning for the second player, i.e., $s_{\omega}(G)=0.$
Conjecture 7.
With the exception of $K_{2}$ and $K_{3}$, for every regular graph $G$, the Star game is winning for the second player, i.e., $s_{\Delta}(G)=0.$
2 Particular behaviour for some variants and extensions
2.1 Smoothening cliques; the games on Colex graphs
Playing the Clique game on $K_{n+1}$ instead of $K_{n}$ implies that there are many more edges to colour. To have a smoother behaviour, we can consider Colex graphs. Let $n$ be the smallest integer for which $\binom{n}{2}\geq m.$ The Colex graphs $G=\mathcal{C}(m)$ of size $m$ can be defined (up to isomorphism) by choosing $V(G)=\{1,2,\ldots,n\}$ and letting $E(G)$ consist of the first $m$ pairs of vertices in colexicographic order (here $(a,b)<(c,d)$ if $b<d$ or $a<c$ and $b=d$). Equivalently, it is the clique $K_{n}$ with the edges of a star of size $\binom{n}{2}-m$ removed.
Figure 1: The Colex graphs $\mathcal{C}(9)$ and $\mathcal{C}(12)$.
A game can be played on any Colex graph $\mathcal{C}(m).$ If the addition of one edge, when going from $\mathcal{C}(m)$ to $\mathcal{C}(m+1),$ would have a small impact in the sense that the outcome $(a,b)$ differs in at most one entry, then by Ramsey theory (or noting the outcomes are not bounded) one could expect infinitely many graphs $\mathcal{C}(m)$ for which the first player can win.
To have even further information, we consider also the Colex variant of the Clique game. Here the players try to form a Colex graph of larger size. Observe that when there is a tie in the Colex game, there is also a tie in the Clique game (and even the exact outcome can be deduced). We verified the outcome of the games for some Colex graphs, the latter summarised in Table 1, and observed that each game exhibits some surprising behaviour.
m $G=\mathcal{C}(m)$ $\operatorname{out}_{\operatorname{col}}(G)$ $\operatorname{out}_{\Delta}(G)$ $s_{\operatorname{VC}}(G)$ $\operatorname{out}_{\operatorname{VC}}(G)$
1 $K_{2}$ (1,0) (1,0) 2 (2,0)
2 $P_{3}$ (1,1) (1,1) 0 (1,1)
3 $K_{3}$ (2,1) (2,1) 1 (1,0)
4 (2,2) (2,2) 1 (2,1)
5 (2,1) (2,1) 2 (2,0)
6 $K_{4}$ (2,2) (2,2) 0 (2,2)
7 (2,2) (3,2) 1 (3,2)
8 (2,2) (3,2) 1 (2,1)
9 $K_{5}^{-}$ (4,2) (3,2) 1 (2,1)
10 $K_{5}$ (4,4) (3,3) 0 (1,1)
11 (4,4) (3,3) 2 (3,1)
12 (4,4) (3,3) 1 (2,1)
13 (4,4) (3,3) 1 (3,2)
14 $K_{6}^{-}$ (4,4) (3,3) 0 (3,3)
15 $K_{6}$ (5,5) (4,4) 0 (3,3)
16 (5,5) (4,4) 1 (4,3)
17 (5,5) (4,3) 1 (3,2)
18 (5,5) (4,4) 0 (3,3)
19 (5,5) (4,4) 1 (3,2)
20 (5,5) (4,4) 0 (2,2)
21 $K_{7}$ (5,5) (4,4) 1 (2,1)
22 (5,5) (4,4) 1 (2,1)
23 (5,5) (4,4) 1 (3,2)
24 (5,5) (4,4) 0 (3,3)
28 $K_{8}$ (5,5) (5,5) 0 (4,4)
Table 1: Outcomes of the three symmetric games on Colex graphs with small size.
  • For the Colex game, the increase of outcomes can occur in at least two ways. Increasing the size from $\mathcal{C}(8)$ to $\mathcal{C}(10),$ implies that initially (for $m=9$) the first player can increase its number of points, but next (at $m=10$) also the second player can do so. Note that $\mathcal{C}(9)$ can be considered as two $K_{4}$s overlapping in a triangle. The first player, Alice, can colour two of the three edges within this triangle and can use this advantage to create a $K_{3}$ in her colour. On the other hand, the extra edge going from $\mathcal{C}(14)$ to $\mathcal{C}(15)$ implies that both players can increase their number of points. The extra edge implies that the playing field is enlarged for both players. We expect this to be the typical behaviour, especially for jumps in outcomes at large sizes.
  • For the Star (maximum degree) game, we observe that Alice wins on $\mathcal{C}(7),\mathcal{C}(8)$ and $\mathcal{C}(9)$. The outcome changes for $16\leq m\leq 18$ When $m=\binom{n-1}{2}+2,$ there is an edge that is the only one in its orbit (e.g. the thick blue one in $\mathcal{C}(12)$ in Fig. 1).
  • In the vertex capturing game, the first player can win on $\mathcal{C}(12)$ despite the order and the size of $\mathcal{C}(12)$ being even.
The behaviour sketched here also suggests that one cannot easily prove that $s_{\omega}(n)=1$ holds in only $o(x)$ many cases for $3\leq n\leq x$.
2.2 The Clique game in less symmetric situations
The question if irregular graphs behave different than cliques was asked in [6, Ques. 1].
Question 8.
Do there exist graphs $H$ with arbitrarily large clique number such that $s_{\omega}(H)=1$?
The main difference between irregular graphs and cliques is that some edges are more important to claim than others in the start, as was the case for $K_{5}^{-}.$ The game played on the graph join $K_{3}+2tK_{1}$ ($K_{3,2t}$ with $3$ edges added between the vertices of degree $2t$) can result in many (t) triangles against none. The latter is true because there are $3$ edges belonging to $12t+$ 1 triangles and the others to only two, so having the advantage to pick two out of three edges with larger impact ($K_{3}$-multiplicity) makes the difference.
If there exists a graph $H$ such that Alice wins the clique game on $H$ whenever a red $K_{3}$ is already present before her move, with outcome $(a,a-1)$ where $a\geq 4,$ then there would be a chance to answer 8 positively for outcome $(a,a-1)$ as well. For this, one could extend $K_{5}^{-}$ by attaching copies $H_{i}$ of the graph $H$, such that $H_{i}\cap K_{5}^{-}\cong K_{3}$ for every $i$, and $V(H_{i})\cap V(H_{j})\subset V(K_{5}^{-})$ for $i\not=j.$
Nevertheless, if Alice plays the clique game on a $K_{8}$, where she starts the game after an advantage of a $K_{3}$ in her colour is already given, the game still ends with $K_{3}$ against $K_{3}$ (outcome $(3,3)$). Starting with a claw, we even observe the following non-smooth jump.
Observation 9.
If Alice starts with a star (claw $K_{1,3}$) $S_{4}$ in her colour as advantage (here the claw spans only maximum degree vertices of the Colex graph) before colouring her first chosen edge, the outcome of the modified Clique game is $(4,4)$ on $K_{8}$ and $(3,3)$ when played on $\mathcal{C}(26)$ or $\mathcal{C}(27).$
In particular, related to 8, we cannot exclude even that for every graph $G$ satisfying $\omega(G)\geq 5$, $\operatorname{Clique}(G)$ is a Player $2$ win.
3 Further results by strategy stealing
In [6], the authors proved that if Alice wins $\operatorname{Clique}(n)$ (something that conjecturally is only the case when $n=2$), then Bob wins $\operatorname{Clique}(n+3)$ for $n\geq R(5,5).$ This was taken over in https://www.erdosproblems.com/778 (at that moment) without the condition of $n$ being sufficiently large. Here we prove that it is indeed a valid statement for all $n$. This follows in essence from [6, Rem. 8] with the additional observation that $n\geq R(3,4)=9$ instead of $n\geq R(4,4)=18$ also works, and the computer check for $n\leq 8$ (checked with independent implementation for $n\leq 7$).
Proposition 10.
If Alice wins $\operatorname{Clique}(n)$ for $n\geq 8$, then Bob wins $\operatorname{Clique}(n+3).$ Bob wins $\operatorname{Clique}(n)$ for $3\leq n\leq 8$
Proof.
The fact that Bob wins $\operatorname{Clique}(n)$ for $3\leq n\leq 8$ was done by the implementation.
Next, we observe that if Alice wins $\operatorname{Clique}(n)$ for $n\geq 8$, she has at least $4$ points, $\omega(G[A])\geq 4$.
Observation 11.
Every graph with order $n=8$ and size $14$ satisfies $\omega(G)\geq 3$ or $\alpha(G)\geq 4$.
Proof.
Easily verified by exhaustive computer check. The three graphs of order $8$ satisfying $\omega(G)=2$ and $\alpha(G)=3$ have size $10,11$ and $12$ (and $R(2,4)<R(3,3)<8$). For more info, one can find these graphs by their HoG Id $588,640,26996$ at https://houseofgraphs.org ([5]). ∎
Now, we show a strategy of Bob to ensure he wins $\operatorname{Clique}(n+3).$ At his first move, he colours an edge adjacent with the edge chosen by Alice. Next, Alice can choose between $5$ configurations up to isomorphism (there are $5$ edge orbits).
Figure 2: The $5$ possible situations after $3$ edges have been coloured.
Bob colours a second edge, not incident with his first, to obtain one of the following three configurations (Fig. 3). Note that his lower blue edge is incident with both of the red edges. One can check each of those $3$ cases separately. Independently of what Alice does in her third turn, Bob can answer, ensuring a triangle $uv_{1}v_{2}$ with two out of three edges coloured blue (called $uv_{1}$ and $uv_{2}$), the subgraph spanned by $V\setminus\{u,v_{1},v_{2}\}$ containing exactly one coloured edge, which is blue. (That is, Bob forces a situation as in [6, Fig. 4] where playing symmetrically, Alice can at most ensure a $K_{4}$ containing $v_{1}v_{2}$.) We show this by giving his action in three cases.
  • If Alice’s third coloured edge is not incident with any of the previous coloured edges, Bob can colour an edge incident to that edge and his lower edge.
  • If Alice chooses an edge which contains one additional vertex, Bob can colour an edge connecting that additional vertex with his lower edge.
  • If Alice chooses an edge among vertices incident to other coloured edges, either Bob can colour an edge incident to a new vertex and his lower edge, or between the lower edge and the unique vertex which was incident to red edges, but none of the blue edges.
By strategy stealing, Bob can obtain a larger clique on the $K_{n}$ spanned by $V\setminus\{u,v_{1},v_{2}\}$ and thus not lose on $K_{n+3}$. For this, note that the largest clique on $K_{n+3}$ coloured by Alice is bounded by the largest one containing $v_{1}v_{2}$ (at most a $K_{4}$) and the largest clique on $V\setminus\{u,v_{1},v_{2}\}$ plus one.
Figure 3: The $3$ possible situations after Bob colours its second edge.
We also present the proof of 5.
Proof of 5..
Let $n\equiv 1\pmod{4}$ where $n\geq 5$. Alice colours one edge, which we call $uv$, red. Bob plays on the other vertices, which induce a $K:=K_{n-2}$, optimally (being the first player there). Whenever Alice colours an edge $ux$ or $vx$, for $x\in V(K),$ Bob colours the other of the two in his colour. If Alice colours an edge in $K,$ so does Bob. Since $s_{\operatorname{VC}}(K)=1,$ and for every $x\in V(K),$ there are two edges outside $K$ with different colours, Bob captures one vertex more among those $n-2$ (by the assumption that  4 is true for $n-2\equiv 3\pmod{4}$). By the pigeon hole principle, one vertex of $u,v$ has at least $\mathopen{}\left\lceil\frac{n-2}{2}\right\rceil\mathclose{}=\frac{n-1}{2}$ incident edges coloured by Bob. Hence Alice will have captured exactly one vertex among $u$ and $v.$ This implies that Bob does not lose. By a stealing argument, Alice will not lose either ([2, Thm. 2.2]), and this implies that $s_{\operatorname{VC}}(K_{n})=0.$
As a corollary of 5, together with the computer verification for $n\leq 8$, we know that 4 is true for $n\leq 9.$
When $4\mid n,$ the truth of 4 looks natural since we know that $s_{\operatorname{VC}}(G)\in\{0,2\}$, while the size and order are both even. Nevertheless, taking into account $\operatorname{VC}(\mathcal{C}(12)),$ a proof is not completely trivial. Also one might think about mirror strategies creating isomorphic coloured graphs on even size cliques $K_{n}.$ Nevertheless, already for $n\in\{4,5\}$ it can be checked that there is no winning strategy for the second player for the creation of isomorphic coloured graphs.
4 Details about computer program
In [3], we included two independent implementations for the computer-aided verifications, one in C and one in Sage. The Sage implementation has simpler syntax, and can be found in the files with name EP&BMI_Games. This section details the strategies used in the C program(s) to solve the games. The publicly available code also serves as a tool for other researchers. We encourage interested readers to use this code as a tool to gain intuition in other variants of the game or the behaviour of the game on different graphs.
4.1 General strategy
The general principle of the program for solving any of the games is divided into two phases. In the first phase, all final configurations are generated. These configurations are independent of the specific game being played and provide the basis for scoring all games. In the second phase, the games are played in reverse. Once we know the outcomes for the (parent) graphs with $r+1$ coloured edges, we consider all children by uncolouring one edge (of the correct colour when playing the game backwards) and remember the optimal score among all parent graphs that lead to this child. This process generates all scores of the graphs with $r$ coloured edges. Proceeding till the graphs with zero coloured edges (or as much as the start configuration), results in the outcome of the game.
4.1.1 Strategies for efficiency in stage one
Apart from the data structures (which we will go into depth in Section 4.1.3) a few techniques are applied to generate the final configurations in an efficient manner. Firstly if the game is played on a clique , we can use a peer-reviewed graph generator from Nauty [7]. The reader can see that, if we only observe final configurations in this game on $n$ vertices, these have $\mathopen{}\left\lceil\frac{n(n-1)}{4}\right\rceil\mathclose{}$ red edges. We can thus generate all non-isomorphic graphs of order $n$ with $\mathopen{}\left\lceil\frac{n(n-1)}{4}\right\rceil\mathclose{}$ edges and assume all edges that are not in this graph to be coloured blue.
This is not the case when either the game is not played on a clique, or player one can start by selecting multiple edges, $E(S)$, on their first turn. For those games a specialised generator was constructed (available in the software package at [3], in the file gen-colex.c). This generator starts from a specific start-graph $S$, representing the edges selected by player one on the first turn, and a base-graph $B$, representing the underlying graph on which the game is played. The assumption here is made that $E(S)\subseteq E(B)$. Then $\mathopen{}\left\lceil\frac{\binom{n}{2}-\lvert E(S)\rvert}{2}\right\rceil\mathclose{}$ edges are placed incrementally, as seen in the function generate_final_configuration. Starting with $S$, we use Nauty to compute the edge-orbits of the graph, the exact procedure to do this correctly will be explained in detail in Section 4.1.4. Then for each orbit only the edge with the lowest index in $S$ is added, as seen in the function compute_possible_edges, and the index of this edge is stored with the graph. The reader can verify this in the function determine_canonical_labelling. Simultaneously we use Nauty to determine the canonical labelling (for details see Section 4.1.4) of this resulting graph and temporarily store this to verify that no isomorphic copies are generated. When all non-isomorphic graphs are generated, the process is repeated on all of these graphs until the required number of edges is reached. An important detail is that by storing the index of the last edge added, we can reduce the amount of isomorphic copies generated. To do this, we allow new edges to be added only with an index higher than the last, ensuring a unique way to add a specific set of edges. This trick is seen in the function compute_possible_edges. This concludes the basic idea of the generator.
4.1.2 Strategies for efficiency in stage two
In the second stage, the game is played in reverse. This phase is performed in erdos-solver.c. To do this, we need all possible final configurations and their scoring (the scoring only needs to adhere to the principle that every game where player one wins has a higher value than every game where player two wins, though more nuanced scoring can also be used), the base-graph $B$ and start-graph $S$. An iterative algorithm similar to the generator is used, seen by the main structure of the function find_best_game. This time however every edge of each colour is removed in turn. In a single iteration we start from a list of non-isomorphic graphs, with sets of edges coloured, either in blue or red, and a set of non-coloured edges. Then for each graph every edge in the colour of the player that is currently playing and that is not in $S$ is removed to generate a graph that represents a possible previous state of the game. Using the techniques described in Section 4.1.4 a canonical representation of the new graph is computed to compare with the graphs already generated in this layer. This is implemented in the function determine_canonical_labelling. Each time a graph is generated that is an isomorphic copy of one that was already generated, only one copy of the graph is kept (with the optimal score). The optimal score is the maximum in the turn of player one and the minimum in the turn of player two. After all possible previous game states have been generated, the process repeats. This iterating stops when only $S$ is found. Then $S$ contains the score that represents the best game player one could play if player two wants to minimise the score of player one.
4.1.3 Datastructures for memory efficiency
To maximise the memory efficiency, the graphs are represented as bitsets. As the edges of the graphs are the most important for our purposes we encode a graph $G$ as follows: label all vertices from 00 to $|V(G)|-1$, then label the edge in colexicographic order; $(1,0)$ as 00, $(2,0)$ as $1$, $(2,1)$ as $2$, $(3,0)$ as $3$, and so on. We now represent each edge with two bits: $00$ represents that the edge is not present in the base-graph and can thus not be coloured, $01$ represents that the edge is not coloured but is present in the base graph and could be coloured, $11$ represents that the edge is present in the base graph, is coloured and has the colour of player one and finally $10$ represents that the edge is present in the base graph, is coloured and has the colour of player two. Thus all information of the graph of order $n$ is encoded in a bitset of $n(n-1)$ bits.
4.1.4 Determining canonical labelling of a coloured graph
To determine the canonical labelling and the edge orbits of a coloured graph we will use Nauty (version 2.8.6) [7]. This library provides a functionality that, given the adjacency matrix of an undirected simple graph with given vertex colour classes, efficiently computes a canonical labelling of the graph and refines the vertex orbits based on the graph structure. The reader can see that the encoding presented in the previous Section 4.1.3 does not comply with this data structure. That is why we expand the graph encoding as follows: Every vertex in the original graph, labelled from 00 to $|V(G)|-1$, is present in the expansion. Then $\frac{n(n-1)}{2}$ new vertices are inserted, each representing a possible edge in the original encoding. These are ordered as described above and labelled from $|V(G)|$ up until $|V(G)|+\frac{n(n-1)}{2}-1$. Lastly four additional vertices are inserted, each representing one of the four possible states of the edge. The vertices representing vertices in the original graph are connected to the vertices representing their respective incident edges in the original graph. The vertices representing edges in the original graph are each connected to the correct vertex representing their state in the original graph encoding. Then to ensure vertices are only mapped to vertices, edge to edges and the states to themselves they are each put into their own colour group. The reader can see this expansion in both the generator and solver in the function generate_expanded_graph. The library is called to canonically label this expanded graph which also returns the vertex orbits for all vertices. We then use the vertex relabelling to construct the bitset of the canonical encoding and store this to compare with other graphs. This is again seen in both implementations in the function determine_canonical_labelling. Additionally, the vertex orbits in the expanded graph of the vertices that represent the edges in the original graph give us the edge-orbits in the original graph. This last part is present in the function compute_possible_edges in the file gen-colex.c.
4.2 Proof of program correctness
In this subsection we explain that the program creates all intermediate colourings and thus exactly executes Zermelo’s theorem (readers knowing the proof of Zermelo’s theorem can skip this).
To prove that the program is correct, we first show that the principle of the generator is sound. To do this, we prove, via induction on the number of red edges in the graph:
It is evident that the starting vertex is the only non-isomorphic copy of itself. Assume we begin with all non-isomorphic graphs with $k$ edges, created using edges with the minimal index. To prove that we then generate all non-isomorphic graphs with $k+1$ edges, created using edges with the minimal index, we assume the contrary: there exists a graph $G$ that is not generated by the algorithm but has $k+1$ edges, all created with edges of minimal index.
Now, there must exist a subgraph $G^{\prime}$ of $G$ that does not contain the edge $e_{h}$ with the highest index in $G$. There are two possibilities for $G^{\prime}$: either it was part of the set of graphs generated at the start of this iteration, or it was not.
In the first case, $G^{\prime}$ would have been extended with each edge in its orbit, using the minimal index. Thus, either the edge was not minimal in its orbit, which contradicts the choice of the edge removed in $G$ to create $G^{\prime}$, or the edge was minimal, but there was another graph isomorphic to $G$ that had already been constructed from another graph $G^{\prime\prime}$ in the starting set. This would contradict the existence of $G$, as the graph generated by adding this other edge to $G^{\prime\prime}$ would have resulted in a graph with edges of a lower index.
If $G^{\prime}$ was not in the starting set, then there is a graph $G^{\prime\prime}$ that is isomorphic to $G^{\prime}$ with edges with a lower index than $G^{\prime}$. Since $G^{\prime\prime}$ has only $k$ edges and is isomorphic to $G^{\prime}$, there exists another edge with smallest index that is not in any edge orbits of edges in $G^{\prime\prime}$ but that is in the edge orbit of one of the edges in $G$. Adding that edge to $G^{\prime\prime}$ would result in a graph isomorphic to $G$ but with edges with lower index, thus contradicting the choice of $G$. This proves no such graph can exist and thus that the strategy results in all non-isomorphic graphs with $k+1$ edges with minimal index.
By induction we proved that the generator generates all non-isomorphic graphs with $\mathopen{}\left\lceil\frac{\binom{n}{2}-\lvert E(S)\rvert}{1+k}\right\rceil\mathclose{}$ edges. Next, we prove that the solver in stage two produces the score of the game if two players play optimally. We prove that the path is optimal by induction on the number of edges to add: we assume that all non-isomorphic final configurations are generated by the generator, and thus each end-configuration is scored as the optimal end-configuration that can be reached from itself.
Next, we assume that we are given all non-isomorphic graphs after $k+1$ moves with the score of the final configuration reached by an optimal game starting from that graph. Assume that the next player is player two, then the current edge was placed by player one. The program uncolours every possible red edge from every graph. If two graphs generated this way are isomorphic, the highest score is stored as the score of the optimal game. This is equivalent to selecting the best possible next graph in each possible graph after $k$ moves. Analogously if the current player is player two, the program uncolours every blue edge and stores the lowest possible score from a given graph after $k$ moves. This is equivalent to selecting the edge that leads to the worst game for player one.
The reader can see that this process converges on the starting graph whereupon the starting graph will have the score of the optimal game.
Acknowledgement
The authors thank Jan Goedgebeur, Jorik Jooken and especially Jarne Renders for early discussions and advice on the efficient implementation of the game.